A company sells 45,000 units of its product at a price of $60 per unit. Variable costs are $51 per unit and fixed costs are $275,000. What is the degree of operating leverage for the firm?

.32

,76

1.47

3.11
Answer #1

Contribution margin=Sales-Variable cost

=(60-51)=$9 per unit

Total contribution margin=(9*45000)=$405,000

Less:Fixed costs=(275,000)

Net operating income=$130,000

Degree of operating leverage=Total contribution margin/Net operating income

=405,000/130,000

=3.11(Approx)